OpenStudy (curry):

2 triangle can be formed using the following info. Solve both. angle B=38, b = 21, c=25

OpenStudy (curry):

i found the big one already. angle c = 47.13 angle a = 94.9 and a =44.6

OpenStudy (anonymous):

using the law of sines we know that \[\frac{\sin(C)}{25}=\frac{\sin(B)}{21}\]

OpenStudy (anonymous):

ok so if you have the big one it should be easy enough to find the small one, but since we may have labeled differently if you post your triangle we can figure out what the other is

OpenStudy (curry):

buti gues my question iss when drawing the 2nd triangle do u disngth of c because hold on 'l sho what i mean

OpenStudy (curry):

disregard the length of c*

OpenStudy (anonymous):

my picture was bad, let me try another one

OpenStudy (anonymous):



OpenStudy (anonymous):

there that is better

OpenStudy (curry):